# Any-stepsize Gradient Descent for Separable Data under Fenchel–Young Losses

### v1

@@ Abstract
The gradient descent (GD) has been one of the most common optimizer in machine learning. In particular, the loss landscape of a neural network is typically sharpened during the initial phase of training, making the training dynamics hover on the edge of stability. This is beyond our standard understanding of GD convergence in the stable regime where arbitrarily chosen stepsize is sufficiently smaller than the edge of stability. Recently, Wu et al. (COLT2024) have showed that GD converges with arbitrary stepsize under linearly separable logistic regression. Although their analysis hinges on the self-bounding property of the logistic loss, which seems to be a cornerstone to establish a modified descent lemma, our pilot study shows that other loss functions without the self-bounding property can make GD converge with arbitrary stepsize. To further understand what property of a loss function matters in GD, we aim to show arbitrary-stepsize GD convergence for a general loss function based on the framework of Fenchel–Young losses. We essentially leverage the classical perceptron argument to derive the convergence rate for achieving $\varepsilon$-optimal loss, which is possible for a majority of Fenchel–Young losses. Among typical loss functions, the Tsallis entropy achieves the GD convergence rate $T=\Omega(\varepsilon^{-1/2})$, and the Rényi entropy achieves the far better rate $T=\Omega(\varepsilon^{-1/3})$. We argue that these better rate is possible because of separation margin of loss functions, instead of the self-bounding property.
@@ 1 Introduction
Gradient-based optimizers are prevalent in the modern machine learning community with deep learning thanks to its scalability and plasticity. Among many variants, the gradient descent remains to be a standard choice. GD with constant stepsize is written as follows:
@@
where $\mathbf{w}\in\mathbb{R}^{d}$ is the optimization variables, $L(\cdot)$ is the loss function, and $\eta>0$ is stepsize fixed across all steps. In convergence analysis of GD, a key ingredient is the descent lemma [36, Section 1.2.3]: for $\beta$-smooth objective $L$, the stepsize choice $\eta<2/\beta$ ensures that $L(\mathbf{w}_{t})$ monotonically decreases. Nonetheless, little optimization theory has been known beyond the threshold $\eta>2/\beta$; though modern neural network exhibits much smaller smoothness values than practically used stepsize values [57, 49]. Moreover, recent studies have reported that GD trajectories of neural networks tend to inflate the sharpness of the loss landscape and hover on the edge of stability (EoS) before convergence [29, 16, 2].
Figure 1: Pilot studies of GD with the same toy dataset as Wu et al. [55]. The dataset consists of four points, $\mathbf{x}_{1}=[1,0.2]^{\top}$, $y_{1}=1$, $\mathbf{x}_{2}=[-2,0.2]^{\top}$, $y_{2}=1$, $\mathbf{x}_{3}=[-1,-0.2]^{\top}$, $y_{3}=-1$, $\mathbf{x}_{4}=[2,-0.2]^{\top}$, and $y_{4}=-1$. GD is run with initialization $\mathbf{w}_{0}=[0,0]^{\top}$. Note that the logistic loss corresponds to the Tsallis $1$-loss. The Tsallis $2$- and $q$-loss are also known as the modified Huber loss [58] and $q$-entmax loss [39], respectively.
Among several recent developments in the theory of arbitrary-stepsize GD (which we will review in Section 1.1), Wu et al. [55] is most closely relevant to ours. They investigated the arbitrary-stepsize behavior of GD by using the binary logistic regression with a linearly separable data, a minimal synthetic setting. They showed that GD initially oscillates with non-monotonic loss values (the EoS phase), which terminates in finite time (phase transition), and then the loss value decreases monotonically (the stable phase). Beyond the logistic loss, these results have been extended to loss functions with the the self-bounding property: for a continuously differentiable loss function $\ell\colon\mathbb{R}\to\mathbb{R}$ and its unsigned derivative $g(\cdot)\coloneqq|\ell^{\prime}(\cdot)|$, $\ell$ satisfies
@@
The self-bounding property generalizes the polynomially-tailed loss introduced by Ji et al. [26] and Ji and Telgarsky [25], and refines the standard smoothness property by allowing the smoothness modulus locally adaptive to the unsigned derivative, such that $C_{\beta}g(x)$. Thus, large $\eta$ can be cancelled out with the vanishingly small loss gradient after the phase transition [55, Lemma 29], and then the GD trajectory follows the descent direction.
In this paper, we study GD with arbitrary stepsize under a wider range of loss functions to identify a key factor to induce the convergent behavior. This is motivated by our pilot study shown in Figure 1, where we found that GD with large stepsize such as $\eta=2^{4}$ remains to converge under the Tsallis $q$-loss (detailed in Section 4), even if the stepsize has gone beyond the classical stable regime. It is noteworthy therein that the Tsallis $q$-loss with $q>1$ does not enjoy the self-bounding property. How much does the self-bounding property play a vital role in arbitrary-stepsize GD convergence? We specifically consider Fenchel–Young losses [9], a class of convex loss functions generated by a potential function $\phi$, as a template of loss functions. Fenchel–Young losses have been used in applications such as structured prediction [37], differentiable programming [8], and model selection [6], while being used as a theoretical tool for online learning [43, 44]. We identify that Fenchel–Young losses with separation margin (formally introduced in Section 2), a relevant notion to the margin in support vector machines, can benefit from better GD convergence rates. Specifically, our main result is informally stated as follows.
Theorem 1 (Informal version of Theorem 5).
Consider a binary classification dataset that is linearly separable. We run (GD) with arbitrary constant stepsize $\eta>0$ and initialization $\mathbf{w}_{0}=\mathbf{0}$ under a Fenchel–Young loss (generated by twice continuously differentiable and convex potential $\phi$) with separation margin. For $\varepsilon>0$, after at most $T$ steps of (GD), where
@@
we have $L(\mathbf{w}_{T})\leq\varepsilon$111 Throughout this paper, we consider $\eta=\Theta(1)$ with respect to the error tolerance $\varepsilon$ when we say arbitrary stepsize unless otherwise noted.
Therefore, (GD) converges regardless of the choice of stepsize $\eta$ under Fenchel–Young losses with separation margin. The order of the convergence rate $T=\Omega(\varepsilon^{-\alpha})$ differs across various potential $\phi$. With a specific choice, the rate can be $T=\Omega(\varepsilon^{-1/2})$ (with $\phi$ being the Tsallis $2$-entropy) and $T=\Omega(\varepsilon^{-1/3})$ (with $\phi$ being the Rényi $2$-entropy, also known as the collision entropy [12]). Remarkably, these convergence rates are better than the classical GD convergence rate $T=\Omega(\varepsilon^{-1})$ under the stable regime, and even better than the convergence rate of the logistic loss after undergoing the EoS and phase transition [55]. Both the Tsallis and Rényi entropies above lack the self-bounding property but entail separation margin. Therefore, we advocate the importance of separation margin for better GD convergence rates. We compare different Fenchel–Young losses in Section 4 and contrast our convergence result with EoS and implicit bias in Section 5.
We present the formal statement of Theorem 1 in Section 3. Its proof significantly leverages the classical perceptron argument [38] without relying on the descent lemma. Intuitively speaking, we track the growth of the parameter alignment $\left\langle{\mathbf{w}_{t}},{\mathbf{w}_{*}}\right\rangle$ with the optimal separator $\mathbf{w}_{*}$. When a loss entails separation margin, $\left\langle{\mathbf{w}_{t}},{\mathbf{w}_{*}}\right\rangle$ cannot grow arbitrarily large (as we simulate in Figure 2 later) while each step of (GD) improves a lower bound of $\left\langle{\mathbf{w}_{t}},{\mathbf{w}_{*}}\right\rangle$, leading to the convergence. Section 3.1 describes this proof overview in detail.
@@ 1.1 Related work
Gradient-based optimization with large stepsize has attracted significant attention recently. Specifically, non-monotonic behaviors of loss functions [56] and the sharpness adaptivity to loss landscapes [29, 16] have been observed empirically. Cohen et al. [16] argued that the sharpness tends to initially increases until the classical stable regime breaks down, and hovers on this boundary, termed as the edge of stability. This observation mainly sparks two questions: why the loss landscape hovers on EoS, and why converging. Answering either question must go beyond the classical optimization theory under the stable regime.
@@
On why the loss landscape hovers on EoS, let us make a brief review, though it is not a central focus of this paper: Ahn et al. [2] is a seminal work to empirically investigate the homogeneity of loss functions contributes to maintain EoS. Later, Lyu et al. [31] showed that normalized GD (represented by scale-invariant losses) adaptively leads their intrinsic stepsize toward sharpness reduction. Ma et al. [32] and Damian et al. [18] attribute the sharpness fluctuation to the non-negligible third-order Taylor remainder of the loss landscape.
@@
We rather focus on why GD converges with arbitrary stepsize. In this line, previous studies show convergence based on specific models such as multi-scale loss function [27], quadratic functions [4], matrix factorization [51, 15], a scalar multiplicative model [59, 28], a sparse coding model [3], and linear logistic regression [54]. Among them, we advocate the logistic regression setup proposed by Wu et al. [54] because it is relevant to implicit bias of GD [47, 24, 40], and moreover, the follow-up work by Wu et al. [55] corroborates the benefit of large stepsize in GD convergence rate. Our work is provoked by Wu et al. [55], questioning what structure in a loss function yields GD convergence. Indeed, we do observe in Figure 1 that loss functions without the self-bounding property (1) can make GD converge, though the self-bounding property seems essential to calm EoS down to the stable phase [55] as well as to establish the max-margin directional convergence [24, 40]. A similar question to ours is raised by Tyurin [50], who argues that the stable convergence of large-stepsize logistic regression might be an artifact due to the functional form of the logistic loss—eventually Tyurin [50] asserts that large-stepsize logistic regression behaves like the classical perceptron. To this end, we show in Theorem 5 that arbitrary-stepsize GD can converge under a wide range of losses even without the self-bounding property (1), and moreover, occasionally yielding a better rate than the classical stable convergence rate. We discuss it more in Section 5. Lastly, we note that Meng et al. [35] attempts to extend the separable logistic regression setup to the non-separable one; yet, we still do not have satisfactory results beyond the one-dimensional case. Due to its intricateness, we follow the linearly separable case for now.
@@ 1.2 Notation
Let $\mathbb{R}_{\geq 0}$ be the set of nonnegative reals. Let $[n]\coloneqq\left\{{1,\dots,n}\right\}$ for $n\in\mathbb{N}$. Let $\mathbf{1}$ be the all-ones vector and $\mathbf{e}_{i}\in\mathbb{R}^{d}$ be the $i$-th standard basis vector, i.e., all zeros except for the $i$-th entry being one. For $\mathcal{S}\subseteq\mathbb{R}^{d}$, $\operatorname{\mathrm{int}}(\mathcal{S})$ denotes its (relative) interior, and $I_{\mathcal{S}}\colon\mathbb{R}\to\left\{{0,\infty}\right\}$ its indicator function, which takes zero if $\bm{\mu}\in\mathcal{S}$ and $\infty$ otherwise. For $\Omega\colon\mathbb{R}^{d}\to\mathbb{R}\cup\left\{{\infty}\right\}$, $\operatorname{\mathrm{dom}}(\Omega)\coloneqq\left\{{\bm{\mu}\in\mathbb{R}^{d}}\mathrel{}\middle|\mathrel{}{\Omega(\bm{\mu})<\infty}\right\}$ denotes its effective domain and $\Omega^{*}(\bm{\theta})\coloneqq\sup\left\{{\left\langle{\bm{\theta}},{\bm{\mu}}\right\rangle-\Omega(\bm{\mu})}\mathrel{}\middle|\mathrel{}{\bm{\mu}\in\mathbb{R}^{d}}\right\}$ its convex conjugate. Let $\triangle^{d}\coloneqq\left\{{\bm{\mu}\in\mathbb{R}_{\geq 0}^{d}}\mathrel{}\middle|\mathrel{}{\left\langle{\mathbf{1}},{\bm{\mu}}\right\rangle=1}\right\}$ be the probability simplex. We introduce $\mathcal{C}^{k}(\mathcal{I})$ as the set of $k$-th continuously differentiable functions on the interval $\mathcal{I}\subseteq\mathbb{R}$.
Let $\Psi\colon\mathbb{R}^{d}\to\mathbb{R}\cup\left\{{\infty}\right\}$ be a strictly convex function differentiable throughout $\operatorname{\mathrm{int}}(\operatorname{\mathrm{dom}}\Psi)\neq\varnothing$. We say $\Psi$ is of Legendre-type if $\lim_{i\to\infty}\|\nabla\Psi(\mathbf{x}_{i})\|_{2}=\infty$ whenever $\mathbf{x}_{1},\mathbf{x}_{2},\dots$ is a sequence in $\operatorname{\mathrm{int}}(\operatorname{\mathrm{dom}}\Psi)$ converging to a boundary point of $\operatorname{\mathrm{int}}(\operatorname{\mathrm{dom}}\Psi)$ (see Rockafellar [42, Section 26]).

### v2

@@ Abstract
The gradient descent (GD) has been one of the most common optimizer in machine learning. In particular, the loss landscape of a neural network is typically sharpened during the initial phase of training, making the training dynamics hover on the edge of stability. This is beyond our standard understanding of GD convergence in the stable regime where stepsize is chosen sufficiently smaller. Recently, Wu et al. [63] have shown that GD converges with much larger stepsize under linearly separable logistic regression. Although their analysis hinges on the self-bounding property of the logistic loss, which seems to be a cornerstone to establish a modified descent lemma, our pilot study shows that other loss functions without the self-bounding property can make GD attain arbitrarily small loss with large stepsize. To further understand what property of a loss function matters in GD, we aim to show large-stepsize GD convergence for a general loss function based on the framework of Fenchel–Young losses. We essentially leverage the classical perceptron argument to derive the iteration complexity for achieving $\varepsilon$-optimal loss, which is possible for a majority of Fenchel–Young losses. This convergence result highlights that the self-bounding property may not be necessary for GD to attain arbitrarily small loss. Moreover, when a loss function entails separation margin, a notion relevant to the margin in support vector machines, GD often yields faster convergence than typical GD rate $T=\Omega(\varepsilon^{-1})$ for convex smooth objectives. Specifically, GD with the Tsallis entropy attains $\varepsilon$-optimal loss with the rate $T=\Omega(\varepsilon^{-1/2})$, and the Rényi entropy achieves the far better rate $T=\Omega(\varepsilon^{-1/3})$.
@@ 1 Introduction
Gradient-based optimizers are prevalent in the modern machine learning community with deep learning thanks to its scalability and plasticity. Among many variants, GD remains to be a standard choice. GD with constant stepsize is written as follows:
@@
where $\mathbf{w}\in\mathbb{R}^{d}$ is the optimization variables, $L(\cdot)$ is the loss function, and $\eta>0$ is stepsize fixed across all steps. The descent lemma [42, Section 1.2.3] is a key to GD convergence: for $\beta$-smooth objective $L$, the stepsize choice $\eta<2/\beta$ ensures that $L(\mathbf{w}_{t})$ monotonically decreases. Nonetheless, little optimization theory has been known beyond the threshold $\eta>2/\beta$; though modern neural networks exhibit much smaller smoothness values than practically used stepsize values [66, 57]. Moreover, recent studies have reported that GD trajectories of neural networks tend to inflate the sharpness of the loss landscape and hover on the edge of stability (EoS) before convergence [34, 18, 2].
Figure 1: Pilot studies of GD with the same toy dataset as [63]. The dataset consists of four points, $\mathbf{x}_{1}=[1,0.2]^{\top}$, $y_{1}=1$, $\mathbf{x}_{2}=[-2,0.2]^{\top}$, $y_{2}=1$, $\mathbf{x}_{3}=[-1,-0.2]^{\top}$, $y_{3}=-1$, $\mathbf{x}_{4}=[2,-0.2]^{\top}$, and $y_{4}=-1$. GD is run with initialization $\mathbf{w}_{0}=[0,0]^{\top}$. Note that the logistic loss corresponds to the Tsallis 11-loss. The Tsallis $2$- and $q$-loss are also known as the modified Huber loss [67] and $q$-entmax loss [45], respectively.
Among several recent developments in the theory of large-stepsize GD (which we will review in Section˜1.1), Wu et al. [63] investigated the large-stepsize behavior of GD by using the binary logistic regression with a linearly separable data, a minimal synthetic setting. They showed that GD initially oscillates with non-monotonic loss values (the EoS phase), which terminates in finite time (phase transition), and then the loss value decreases monotonically (the stable phase). Beyond the logistic loss, these results have been extended to loss functions with the self-bounding property: for a differentiable loss function $\ell\colon\mathbb{R}\to\mathbb{R}$ and its absolute derivative $g(\cdot)\coloneqq|\ell^{\prime}(\cdot)|$, $\ell$ satisfies
@@
The self-bounding property generalizes the polynomially-tailed loss [31, 30], and refines the standard smoothness property by allowing the smoothness modulus locally adaptive to the derivative, such that $C_{\beta}g(x)$. Thus, large $\eta$ can be cancelled out with the vanishingly small loss gradient after the phase transition [63, Lemma 29], and GD follows the descent direction.
In this paper, we study GD with large stepsize under a wide range of loss functions to identify a key factor to induce the convergent behavior. This is motivated by our pilot study shown in Figure˜1, where we found that GD with large stepsize such as $\eta=2^{4}$ remains to converge under the Tsallis $q$-loss (detailed in Section˜4), even if the stepsize has gone beyond the classical stable regime. It is noteworthy therein that the Tsallis $q$-loss with $q>1$ does not enjoy the self-bounding property. How much does the self-bounding property play a vital role in large-stepsize GD convergence? We specifically consider Fenchel–Young losses [11], a class of convex loss functions generated by a potential function $\phi$, as a template of loss functions. Fenchel–Young losses have been used in applications such as structured prediction [43], differentiable programming [10], and model selection [7], while being used as a theoretical tool for online learning [51, 52]. We identify that Fenchel–Young losses with separation margin (formally introduced in Section˜2), a relevant notion to the margin in support vector machines, can often benefit from better GD convergence rates. We say a loss function has separation margin if the loss value vanishes with a sufficiently large positive prediction margin. Specifically, our main result is informally stated as follows.
Theorem 1 (Informal version of Theorem˜5).
Consider a binary classification dataset that is linearly separable. We run (GD) with arbitrary constant stepsize $\eta>0$ and initialization $\mathbf{w}_{0}=\mathbf{0}$ under a Fenchel–Young loss generated by twice continuously differentiable and convex potential $\phi$ with separation margin. For $\varepsilon>0$, after at most $T$ steps of (GD), where
@@
we have $L(\mathbf{w}_{T})\leq\varepsilon$. [^1] Throughout this paper, we consider $\eta=\Theta(1)$ with respect to the error tolerance $\varepsilon$ when we say arbitrary stepsize unless otherwise noted.
As defined in Section˜2, a loss function with separation margin vanishes for a sufficiently large prediction margin, which is a natural indicator of correct classification used in support vector machines. The order of the convergence rate $T=\Omega(\varepsilon^{-\alpha})$ differs across various potential $\phi$. With a specific choice, the rate can be $T=\Omega(\varepsilon^{-1/2})$ (with $\phi$ being the Tsallis $2$-entropy) and $T=\Omega(\varepsilon^{-1/3})$ (with $\phi$ being the Rényi $2$-entropy, also known as the collision entropy [14]). Remarkably, these convergence rates are better than the classical GD convergence rate $T=\Omega(\varepsilon^{-1})$ under the stable regime, and even better than the convergence rate of the logistic loss after undergoing the EoS and phase transition [63]. Both the Tsallis and Rényi entropies above lack the self-bounding property but have separation margin. Therefore, we advocate the importance of separation margin for better GD convergence rates. We compare different Fenchel–Young losses in Section˜4 and contrast our convergence result with the EoS and implicit bias in Section˜5.
We present Theorem˜1 formally in Section˜3. Our proof leverages the classical perceptron argument [44] without relying on the descent lemma at all. Intuitively speaking, we track the growth of the parameter alignment $\left\langle{\mathbf{w}_{t}},{\mathbf{w}_{*}}\right\rangle$ with the optimal separator $\mathbf{w}_{*}$. When a loss entails separation margin, $\left\langle{\mathbf{w}_{t}},{\mathbf{w}_{*}}\right\rangle$ cannot grow arbitrarily large (as we simulate in Figure˜2 later) while each step of (GD) improves a lower bound on $\left\langle{\mathbf{w}_{t}},{\mathbf{w}_{*}}\right\rangle$, leading to the convergence. Section˜3.1 describes this proof overview in detail. This is different from the proof of Wu et al. [63], whose core is the modified decent lemma (recapped in Lemma˜21 in the appendix) based on the self-bounding property. Although the perceptron argument is partially used therein [63], the average loss is finally controlled by the modified descent lemma, and thus the proof is only applicable to the self-bounding losses.
@@ 1.1 Related work
Gradient descent with large stepsize has attracted attention recently. Specifically, non-monotonic behaviors of loss functions [65] and the sharpness adaptivity to loss landscapes [34, 18] have been observed empirically. It was argued that the sharpness tends to initially increases until the classical stable regime breaks down, and hovers on this boundary, termed as the edge of stability [18]. This observation mainly sparks two questions: why the loss landscape hovers on the EoS, and why converging. Answering either question must go beyond the classical optimization theory under the stable regime.
@@
On why hovering on the EoS, let us make a brief review, though it is not a central focus of this paper: Ahn et al. [2] is a seminal work to empirically investigate the homogeneity of loss functions contributes to maintain the EoS. Later, it was showed that normalized GD (represented by scale-invariant losses) adaptively leads their intrinsic stepsize toward sharpness reduction [36]. The sharpness fluctuation is often attributed to the non-negligible third-order Taylor remainder of the loss landscape [37, 20].
@@
We rather focus on why GD attains arbitrarily small loss with much larger stepsize. In this line, previous studies show convergence based on specific models such as multi-scale loss function [32], quadratic functions [5], matrix factorization [59, 17], a scalar multiplicative model [68, 33], a sparse coding model [3], and linear logistic regression [62]. Among them, we advocate the logistic regression setup proposed by Wu et al. [62] because it is relevant to implicit bias of GD [54, 29, 46], and moreover, Wu et al. [63] corroborates the benefit of large stepsize in GD convergence rate. Our work is provoked by Wu et al. [63], questioning what structure in a loss function leads GD to arbitrarily small loss. Indeed, we observe in Figure˜1 that loss functions without the self-bounding property (1) can make GD attain arbitrarily small loss, though the self-bounding property seems essential to calm the EoS down to the stable phase [63] as well as to establish the max-margin directional convergence [29, 46]. A similar question to ours is raised by Tyurin [58], who argues that the stable convergence of large-stepsize logistic regression might be an artifact due to the functional form of the logistic loss—eventually Tyurin [58] argued that large-stepsize logistic regression behaves like the classical perceptron. To this end, we show in Theorem˜5 that arbitrary-stepsize GD can converge under a wide range of losses even without the self-bounding property (1), and moreover, occasionally yielding a better rate than the classical stable convergence rate. We discuss it more in Section˜5. Note that one work attempts to extend the separable logistic regression setup to the non-separable one [41]; yet, we still do not have satisfactory results beyond the one-dimensional case. Due to its intricateness, we follow the separable case.
Lastly, our work benefits the study of regret bounds of surrogate losses [9, 1, 23, 6, 38, 8]. A surrogate regret bound connect a surrogate loss to a downstream task loss, while the optimization error of the surrogate loss is usually ignored. Our GD convergence analysis can be integrated to surrogate regret bounds when discussing a downstream task performance.
@@ 1.2 Notation
Let $\mathbb{R}_{\geq 0}$ be the set of nonnegative reals. Let $n]\coloneqq\left\{{1,\dots,n}\right\}$ for $n\in\mathbb{N}$. Let $\mathbf{1}$ be the all-ones vector and $\mathbf{e}_{i}\in\mathbb{R}^{d}$ be the $i$-th standard basis vector, i.e., all zeros except for the $i$-th entry being one. For $\mathcal{S}\subseteq\mathbb{R}^{d}$, $\operatorname{\mathrm{int}}(\mathcal{S})$ denotes its (relative) interior, and $I_{\mathcal{S}}\colon\mathbb{R}\to\left\{{0,\infty}\right\}$ its indicator function, which takes zero if $\bm{\mu}\in\mathcal{S}$ and $\infty$ otherwise. For $\Omega\colon\mathbb{R}^{d}\to\mathbb{R}\cup\left\{{\infty}\right\}$, $\operatorname{\mathrm{dom}}(\Omega)\coloneqq\left\{{\bm{\mu}\in\mathbb{R}^{d}}\mathrel{}\middle|\mathrel{}{\Omega(\bm{\mu})<\infty}\right\}$ denotes its effective domain and $\Omega^{*}(\bm{\theta})\coloneqq\sup\left\{{\left\langle{\bm{\theta}},{\bm{\mu}}\right\rangle-\Omega(\bm{\mu})}\mathrel{}\middle|\mathrel{}{\bm{\mu}\in\mathbb{R}^{d}}\right\}$ its convex conjugate. Let $\triangle^{d}\coloneqq\left\{{\bm{\mu}\in\mathbb{R}_{\geq 0}^{d}}\mathrel{}\middle|\mathrel{}{\left\langle{\mathbf{1}},{\bm{\mu}}\right\rangle=1}\right\}$ be the probability simplex. We introduce $\mathcal{C}^{k}(\mathcal{I})$ as the set of $k$-th continuously differentiable functions on the interval $\mathcal{I}\subseteq\mathbb{R}$.
Let $\Psi\colon\mathbb{R}^{d}\to\mathbb{R}\cup\left\{{\infty}\right\}$ be a strictly convex function differentiable throughout $\operatorname{\mathrm{int}}(\operatorname{\mathrm{dom}}\Psi)\neq\varnothing$. We say $\Psi$ is of Legendre-type if $\lim_{i\to\infty}\|\nabla\Psi(\mathbf{x}_{i})\|_{2}=\infty$ whenever $\mathbf{x}_{1},\mathbf{x}_{2},\dots$ is a sequence in $\operatorname{\mathrm{int}}(\operatorname{\mathrm{dom}}\Psi)$ converging to a boundary point of $\operatorname{\mathrm{int}}(\operatorname{\mathrm{dom}}\Psi)$ (see [49, Section 26]).